For the total cos function gives as TC=15+3Q-1.5Q2+7.5Q3-45Q4. Compute the Tfc, Tvc, Afc, Avc and Atc
1
Expert's answer
2021-08-09T06:47:43-0400

Solution:

i.). The Total Fixed Cost (TFC) = TC – TVC = (15 + 3Q – 1.5Q2 + 7.5Q3 – 45Q4) – (3Q – 1.5Q2 + 7.5Q3 – 45Q4) = 15

The Total Fixed Cost (TFC) = 15

 

ii.). The Total Variable Cost (TVC) = 3Q – 1.5Q2 + 7.5Q3 – 45Q4

 

iii.). Average Fixed Cost (AFC) = "\frac{TFC}{Q}" = "\frac{15}{Q}"


Average Fixed Cost (AFC) = "\frac{15}{Q}"

 

iv.). Average Variable Cost (AFC) = "\frac{TVC}{Q} = \frac{3Q - 1.5Q^{2} + 7.5Q^{3} - 45Q^{4} }{Q} = 3 - 1.5Q + 7.5Q^{2} - 45Q^{3}" 3Q – 1.5Q2 + 7.5Q3 – 45Q4/Q = 3 – 1.5Q + 7.5Q2 – 45Q3

Average Variable Cost (AFC) = "3 - 1.5Q + 7.5Q^{2} - 45Q^{3}"

 

v.). Average Total Cost (ATC) = "\frac{TC}{Q} = \frac{15 + 3Q - 1.5Q^{2} + 7.5Q^{3} - 45Q^{4} }{Q} = \frac{15}{Q}+3 - 1.5Q + 7.5Q^{2} - 45Q^{3}"


Average Total Cost (ATC) = "\frac{15}{Q}+3 - 1.5Q + 7.5Q^{2} - 45Q^{3}"